A manufacturer is trying to determine the combination of two products, A and B, that should be produced. Each product passes through a three-stage production process: Stage I, II, III. For product A Stage I takes three hours of labour, Stage II four hours and Stage III one hour. For product B the comparable data is four hours, two hours and two hours. Labour hours currently cost £4.50 per hour. Product A sells for £38 a unit and B for £40. At present 11 people are employed on Stage I, 12 on Stage II and 10 on Stage III, with all employees working 40 hours each week.

The manufacturer depicted above initially decides to find a product mix that maximizes revenue. What will be the combination of products produced that maximizes revenue? A second study is undertaken to find the mix of products that maximizes profits. Using solver on Excel, find the combination of products A and B to be produced that will maximize profits. What is the significant difference in these results? Why are they different? Which approach would you recommend? Explain your response.
Answer #1

Total labour hrs required for product in all stages A= 3+4+1 = 8 Hrs

Total labour hrs required for product in all stages B= 4+2+2 = 8 Hrs

A B C F 1 2. Labour Total Labour Hrs 3 Stage 1 4 Stage 2 5 Stage 3 11 12 10 440 480 400 D E Labour Hrs utilised Product A ProE13 X fi с D F Total Labour Hrs =B3*40 =B4*40 =B5*40 E Labour Hrs utilised Product A Productb B =$D$9*B9 =$E$9*C9 =$D$9*B10 |Solver Parameters X Set Objective: SCS18 To: Max Min O Value Of: 0 By Changing Variable Cells: SD$9:$E$9 17 Subject to the CoA B с F 1 Total Labour Hrs 440 480 400 E Labour Hrs utilised Product A Productb B 0 440 0 220 0 220 0 880 440 220 220 Product